OpenStudy (anonymous):

Given that (2-3i)a + 3b = 2 + 5i, find the values of a and b such that a and b are conjugate complex numbers. I know how to solve this for a and b are both real, how would I solve it for complex conjugates? Thanks.

OpenStudy (anonymous):

oh wait that is wrong.

OpenStudy (anonymous):

Let a=(x+iy) and b=(x-iy)

OpenStudy (anonymous):

Now substitute them to find x and y

OpenStudy (anonymous):

let my multiply correctly oh no, it is right! \[2a-3ai+3b=2+5i\] \[(2a+3b)-3ai=2+5i\]

OpenStudy (anonymous):

oooooooh i see. a and be themselves are complex numbers. sorry. roshan is right

OpenStudy (anonymous):

Oh, right. Thanks.

OpenStudy (anonymous):

The answers are gross >.>

OpenStudy (zarkon):

yep